A muon produced in the Earth’s atmosphere is travelling with a speed of 0.90c. As

measured in the muon’s frame of reference, it has a lifetime of 1.6 us. What is its

lifetime as measured by an observer on Earth?


1
Expert's answer
2021-03-23T11:10:46-0400

The lifetime as measured by an observer on Earth can be found using the relativistic time dilation (see https://en.wikipedia.org/wiki/Time_dilation#Simple_inference):


"\Delta t' = \dfrac{\Delta t}{\sqrt{1-v^2/c^2}}"

where "\Delta t'" is the lifetime as measured by an obsever, "\Delta t = 1.6\mu s = 1.6\times 10^{-6}s" is the lifetime in the frame of the muon, "v = 0.9c" is the speed of the muon, and "c" is the speed of light. Thus, obtain:


"\Delta t' = \dfrac{1.6\times 10^{-6}}{\sqrt{1-0.9^2c^2/c^2}} \approx 3.67\times 10^{-6}s"

Answer. 3.67 us.